OpenStudy (anonymous):

2/5x + 2/5(x-10) = 16

OpenStudy (daggiemib):

https://mathway.com/

OpenStudy (anonymous):

That is such a cheat site!

OpenStudy (anonymous):

YOU HAVE TO PAY FOR THE STEPS!

OpenStudy (daggiemib):

PEMDAS 2/5+2/5(x-10)=16 distribute 2/5 2/5+ 2/5x -4=16 combine like terms 2/5x-3.6=16 add 3.6 on both sides 2/5x=19.6 divide both sides by 2/5 x=49

OpenStudy (daggiemib):

PRACTICE PEMDAS AND YOU WILL BE A PRO

OpenStudy (anonymous):

THX SOOOO MUCH!!

OpenStudy (daggiemib):

oops

OpenStudy (anonymous):

but hey we all make mistakes but you showed my the right path

OpenStudy (daggiemib):

then add like terms and youll get it!:) no problem